keiko needs 100ml of a 5% acid solution for a science experiment. She has available a 1% solution and how many millikiters of the 6% solution should she mix to make the 5% solution
Answer #1

: Percent Quantity 1% - de 6% - 100.de 5% - 100 According to conditions - 18+ 6( 100-4e) : 51100) H + 600 -6 H = 500 -54 = 50